OpenStudy (anonymous):

Evaluate sin(Cos^-1 * (-15/17)) .

OpenStudy (anonymous):

Ok ... what sides of a right triangle are the ratio for cos?

OpenStudy (anonymous):

So do you just want someone to give you the answer or do you want help?

OpenStudy (anonymous):

It seems that you are just fishing for an answer. Good luck.

OpenStudy (anonymous):

the answer is 0.470588235

OpenStudy (anonymous):

Sorry, I was away from my computer @Paynesdad!

OpenStudy (anonymous):

Not exactly. But that is the approximate answer.

OpenStudy (anonymous):

I want to know how to get that answer. Besides I need a fraction, lolz

OpenStudy (anonymous):

cos = adjacent/hypotenuse

OpenStudy (mathstudent55):



OpenStudy (anonymous):

\[\cos (\Theta) = \frac{-15}{17} = \cos^{-1}(\frac{-15}{17})\] so draw a triangle  Use the Pythagorean solve for opposite and \[\sin (\Theta) = \frac{opposite}{Hypotenuse}\]

OpenStudy (anonymous):

OH, nice visual

OpenStudy (mathstudent55):

@ilfy214 Do you understand my figure above? You can get your answer directly from it.

OpenStudy (anonymous):

-15/17

OpenStudy (mathstudent55):

Your problem has two parts. 1. What is arccos of (-15/17)? 2. What is sin of angle in part 1?

OpenStudy (anonymous):

I'm sorry. that was cos. 8/17

OpenStudy (mathstudent55):

Look in my figure. I drew the side of 15 as -15 in the graph so you see the angle whose cosine in -15/17.

OpenStudy (mathstudent55):

Now take the sine to get 8/15. I see that you got it. Great!

OpenStudy (anonymous):

Thanks! But you mean 8/17, right? :)

OpenStudy (mathstudent55):

yes, 8/17

OpenStudy (anonymous):

Nice job!

OpenStudy (anonymous):

THANK YOU EVERYONE!